# Descriptive Caption Enhancement with Visual Specialists for Multimodal Perception

### v2

@@ Abstract
Training Large Multimodality Models (LMMs) relies on descriptive image caption that connects image and language. Existing methods either distill the caption from the LMM models or construct the captions from the internet images or by human. We propose to leverage off-the-shelf visual specialists, which were trained from annotated images initially not for image captioning, for enhancing the image caption.
Our approach, named DCE, explores object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories) and object relations (e.g., relative location and human-object-interaction (HOI)), and combine the attributes into the descriptive caption. Experiments demonstrate that such visual specialists are able to improve the performance for visual understanding tasks as well as reasoning that benefits from more accurate visual understanding. We will release the source code and the pipeline so that other visual specialists are easily combined into the pipeline. The complete source code of DCE pipeline and datasets will be available at https://github.com/syp2ysy/DCE.
@@ 1 Introduction
Recent advancements in Large multimodal models (LMMs) [53, 55, 60] have significantly enhanced the understanding and reasoning abilities for multimodal tasks. Vision-language connection is crucial for high abilities, and descriptive image captions serve as one of key components for image perception. The image caption is expected to describe the image as detailed and complete as possible. There are two main categories for image captions. One is to generate image captions from human annotation, such as COCO [30]and LAION [43]. However, the high cost of human annotation limits scalability. The other one is LMMs [11, 1] annotation, such as ShareGPT4V [8] and Densefusion [26]. While captions generated by LMMs offer better scalability, their comprehensiveness and accuracy often fall short.
There is still much improvement space for the captions from SoTA LMMs, such as InternVL2 [11] and from current human annotation datasets. An example is given in Figure 1(a) for showing the improvement space. It can be seen that the caption from COCO [29] a widely-used human-annotated dataset, is usually incomplete, only describes a small portion of the image content, and a lot of information is missing. The captions from LMMs are much better, but still can be improved in some aspects. As shown in Figure 1(b), we selected 8 objects and 5 key attributes from the image to analyze and compare the captions generated by different methods. The key attributes include fine-grained attributes, spatial relation, and HOI. It is clear that captions generated by LMMs are more detailed than those annotated by humans. They not only describe more objects but also more attributes. However, captions from LMMs still much improvement space, as they tend to overlook important objects and attributes. For example, in Figure 1(a), Object 6 is completely ignored by all LMMs. Furthermore, crucial 3D spatial relationships between objects are also missing, which is particularly problematic for tasks that require a comprehensive understanding of the scene’s structure [21, 34].
@@
Figure 1: (a) We present a comparison of captions from DCE, human, and generalist LMM models annotations, including InternVL2-26B, LLaVA-NeXT, and GPT-4V. (b) visualizes the extent to which the captions in (a) describe multiple objects and various attributes, including Objects 1-8, Object Attributes, OCR, HOI, 2D spatial relations and 3D spatial relations.
Toward this end, we propose a Descriptive Caption Enhancement Engine (DCE), designed to enable efficient and low cost image captioning. We leverage visual specialists [59, 9, 48, 54] to replicate various human visual capabilities, and subsequently employ large language models (LLMs) [51, 4] to simulate the human cognitive process. This combined approach enables us to generate high-quality image captions by closely mimicking the way humans perceive and interpret visual information. Notably, DCE relies solely on open-source visual expert models and LLMs, significantly reducing annotation costs.
Specifically, we leverage existing visual specialists to obtain instance-level and relational attributes within images. Instance-level attributes focus on object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories). Relational-level attributes capture interactions and relationships between objects (e.g., relative location and HoI). Next, we use prompts to guide LLMs in combining object attributes into region captions. Finally, prompts are used again to integrate these region captions with relational-level attributes, producing a comprehensive and detailed image caption. Since DCE utilizes multiple off-the-shelf visual specialists, the resulting captions capture a wide array of detailed attributes and nuanced relationships, leading to richer and more precise image descriptions. As shown in Figure 1(b), captions annotated by DCE contain the most comprehensive information among all methods. Figure 2 quantitatively demonstrates that the captions generated by DCE provide greatest benefit to LMMs.
We applied our DCE to annotate a large-scale dataset of 1.1 million images, consisting of 1 million diverse images (DCE-1M) and 118K real-world scene images (DCE-118K). Experiments were conducted using both LLaVA-v1.5 and LLaVA-NeXT models. The results show that the highly detailed captions generated by DCE significantly enhance the perceptual capabilities of large multimodal models (LMMs), improving visual-language alignment. The outstanding performance of both LLaVA-v1.5 and LLaVA-NeXT across 14 benchmarks further underscores the effectiveness of our approach, confirming that the generated image captions are of exceptional detail and quality.
@@ 2 Related Work
Figure 3: The DCE pipeline first utilizes various visual specialists to extract both Object and Relation attributes. Then, it uses an LLM to integrate the object attributes into detailed region captions, followed by combining the region captions with relational attributes to generate a comprehensive image caption.
@@
Depth Depth & Detection model Average the depth map values within the bounding box region to obtain the depth information.
Emotion Emotion model If the detected region is labeled as ”person”, an emotion model is used to extract an emotion label.
OCR OCR Model Using an OCR model to extract the text content and bounding box from the region.
Animal Fine Grained model A fine-grained recognition model to identify specific species of the animal.
Plants A fine-grained recognition model to identify specific species of the plants.
Aircrafts A fine-grained recognition model to identify specific model of the aircraft.
Logo A fine-grained recognition model to identify logos in the region.
Landmark A fine-grained recognition model to identify landmarks within the region.
Food A fine-grained recognition model to identify specific species of the food.
Celebrity Using a fine-grained recognition model to identify celebrity within the region.
@@
Count Detection model Counting the number of all objects in the image based on the detection results.
2D Absolute Location Detection model
@@
2D Relative Location Detection model
@@
3D Relative Location Detection & Depth model
@@
Large multi-modality models. There are recently a lot of developments for large multi-modality models. One major effort lies in aligning/connecting the pretrained vision encoder and the the pretrained language model [53, 55, 50]. Flamingo [2] inserts new gated cross-attention layers between existing pretrained and frozen LM layers, that bridges powerful pretrained vision-only and language-only models. BLIP-2 [25] bridges the two modalities with a lightweight querying transformer. Qwen-VL [4] adopt a way that is similar to query transformer. LLaVA [33] adopt a simple projection model to connect the visual encoder and the language model. The Llama3 Herd of Models [14] uses a adapter similar to LLaVA and BLIP-2 to connect the vision encoder and the language model. Gemini [16] simply concatenates multi-modality tokens, such as image and text tokens, and feed them into a transformer as the input. Emu [49] uses a causal transformer to train the image encodings to tokens. There is also some effort lying in the encoder architecture for efficient image encoding. One approach involves dividing the image into large blocks, which helps capture finer-grained features while maintaining computational efficiency, such as Monkey [28], Qwen2-VL [52], LLaVA-NeXT [24]. Additionally, many works have further enhanced LMMs by focusing on the quality and diversity of pretraining and finetuning data [4, 40, 22, 7, 6, 53]. Despite the success of these LMMs, few studies have focused on obtaining large-scale high-quality image-text data, which is a crucial factor in driving the capabilities of LMMs.
Descriptive captions. CC3M [45] (Conceptual Captions: A Cleaned, Hypernymed, Image Alt-text Dataset For Automatic Image Captioning) is harvested from the Alt-text HTML attribute associated with web images followed by an automatic pipeline that extracts, filters, and transforms candidate image/caption pairs. CC12M [5] extends CC3M [45] with an emphasis on long-tail visual recognition. It builds on the same foundational methodology as CC3M but incorporates a more diverse range of objects, scenes, and rare visual concepts to better represent the breadth of real-world imagery. SBU Captions [39] forms the data with images and descriptions sourced from Flickr. The captions are manually written by users and describe a wide variety of visual content, from everyday scenes to more specialized imagery. COCO-captions [10] is built by collecting captions on Flickr images that are generated by human subjects on Amazon’s Mechanical Turk (AMT). While manually annotated captions effectively capture the main information in images, they often overlook important details and contextual richness. To address this, methods like ShareGPT4V [8] and Densefusion [26] leverage LLMs [8, 52] to generate more detailed captions. However, these LMM-based approaches still face challenges, particularly with over-simplified or inaccurate descriptions, which can compromise caption quality and reliability [41, 21]. Balancing detail richness and accuracy remains a key challenge in current research. DenseFusion and our DCE share a similar goal of generating more accurate image captions by integrating diverse visual information. However, DenseFusion relies on GPT-4V, a highly expensive model, whereas our approach significantly reduces costs and improves efficiency by utilizing open-source visual expert models and LLMs.
@@ 3 Approach
As shown in Figure 3, DCE leverage existing visual specialists to extract visual properties for improving descriptive captions. We explore two kinds of properties: object-level attributes and object-relations. Our approach consists of: object localization using a SoTA open-world object detection specialist, object property extraction using various specialists, and object relation extraction between objects.
@@ 3.1 Object Attributes
Attribute extraction by visual specialists. Table 1 presents all the instance-level attributes involved in DCE, along with their extraction processes. The attributes currently include three main parts. (1) Fine-grained object category attributes. This is rarely explored in current multimodal models [8]. In DCE, we incorporate fine-grained details by introducing various specialized models, covering categories such as animals, plants, food, logo, aircraft, landmarks, and celebrities. The animal and plant attributes contain the fine-grained category of $891k$ and $427k$ species of animals and plants, respectively. Each species within this category possesses unique characteristics and behaviors, making it a rich and varied classification. The food attributes include a variety of food types commonly found in daily life, while logos are graphic symbols that serve as visual identifiers for conveying messages. Landmarks are the famous tourist attractions that hold cultural, historical, or geographical significance. Aircraft refers to the machines designed for flight, including airplanes, helicopters, drones, and other aerial vehicles. Celebrities are individuals widely recognized in public life, entertainment, sports, or other fields. All this specific and fine-grained information could be regarded as the external world knowledge, aligning the textual content in the image with basic human cognition. (2) Low-level and emotion attributes. It includes emotion, depth, and size. (3) OCR. It is one of the important attributes for multimodal models. The specific visual expert models we use will be presented in the Appendix.
@@ 3.2 Object Relation
DCE can extract relationships between multiple objects. We consider three categories: the interactions between humans and objects, the 2D as well as 3D relative positional relationships among different objects, and the object counting information. Table 1 presents the relation attributes and their extraction process.
@@
Attributes InternVL2 LLaVA-NeXT DCE

### v3

# Enhancing Descriptive Captions with Visual Attributes for Multimodal Perception

@@ Abstract
Training Large Multimodality Models (LMMs) relies on descriptive image caption that connects image and language. Existing methods for generating such captions often rely on distilling the captions from pretrained LMMs, constructing them from publicly available internet images, or even generating them through human annotation. However, these strategies can fall short in terms of precision and granularity, particularly when dealing with complex visual reasoning tasks. In this paper, we propose to leverage off-the-shelf visual specialists, which were trained from annotated images initially not for image captioning, for enhancing the image caption. Our approach, named EDC, explores object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories) and object relations (e.g., relative location and human-object-interaction (HOI)), and combine the attributes into the descriptive caption. By systematically integrating these rich attributes into the generated captions, EDC significantly improves the descriptive quality of the captions, providing a deeper and more nuanced understanding of the visual content. Experiments demonstrate that such visual specialists are able to improve the performance for visual understanding tasks as well as reasoning that benefits from more accurate visual understanding.
@@ 1 Introduction
Recent advancements in Large multimodal models (LMMs) [64, 77] have significantly enhanced the understanding and reasoning abilities for multimodal tasks. Vision-language connection is crucial for high abilities, and descriptive image captions serve as one of key components for image perception [68, 19]. The image caption is expected to describe the image as detailed and complete as possible. There are two main categories for image captions. One is to generate image captions from human annotation [14, 72], such as COCO [33]and LAION [50]. However, the high cost of human annotation limits scalability. The other one is LMMs [1] annotation, such as ShareGPT4V [7] and Densefusion [29]. While captions generated by LMMs offer better scalability, their comprehensiveness and accuracy often fall short.
There is still much improvement space for the captions from SoTA LMMs, such as InternVL2 [10] and from current human annotation datasets, which we illustrate through concrete examples in Figure 1. First, human-annotated captions tend to be incomplete, often describing only the most salient objects while omitting finer-grained details and contextual relationships. For example, in Figure 1(a), the widely used COCO dataset provides human annotations that capture only a part of the scene. These captions typically focus on high-level descriptions (e.g., “a shirtless man with a green tattoo is skateboarding”) while ignoring crucial elements such as the number of objects, their fine-grained attributes, or spatial relationships. This oversimplification limits the utility of such captions, particularly in enhancing the vision-language alignment of LMMs, as it restricts their ability to develop a comprehensive understanding of image content.
@@
Figure 1: (a) We present a comparison of captions from EDC, human, and generalist LMM models annotations, including InternVL2-26B, LLaVA-NeXT, and GPT-4V. (b) visualizes the extent to which the captions in (a) describe multiple objects and various attributes, including Objects 1-8, Object Attributes, OCR, HOI, 2D spatial relations and 3D spatial relations.
@@
On the other hand, captions generated by SoTA LMMs [1, 10], while an improvement over human annotations, still exhibit notable deficiencies. As shown in Figure 1 (b), we selected 8 objects and 5 key attributes from the image to analyze and compare the captions generated by different methods. The key attributes include fine-grained attributes, spatial relation, and HOI. It is clear that captions generated by LMMs are more detailed than those annotated by humans. They not only describe more objects but also more attributes. However, captions from LMMs still much improvement space, as they tend to overlook important objects and attributes. For example, in Figure 1(a), Object 6 is completely ignored by all LMMs. Furthermore, crucial 3D spatial relationships between objects are also missing, which is particularly problematic for tasks that require a comprehensive understanding of the scene’s structure [24].
Toward this end, we propose an Enhancing Descriptive Captions Engine (EDC), designed to enable efficient and cost-effective image captioning. We leverage visual specialists [76, 8, 57, 65] to replicate various human visual capabilities, and subsequently employ large language models (LLMs) [60] to simulate the human cognitive process. This combined approach enables us to generate high-quality image captions by closely mimicking the way humans perceive and interpret visual information. Notably, EDC relies solely on open-source visual expert models and LLMs, significantly reducing annotation costs.
Specifically, we leverage existing visual specialists to obtain instance-level and relational attributes within images. Instance-level attributes focus on object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories). Relational-level attributes capture interactions and relationships between objects (e.g., relative location and HoI). Next, we use prompts to guide LLMs in combining object attributes into region captions. Finally, prompts are used again to integrate these region captions with relational-level attributes, producing a comprehensive and detailed image caption. Since EDC utilizes multiple off-the-shelf visual specialists, the resulting captions capture a wide array of detailed attributes and nuanced relationships, leading to richer and more precise image descriptions. As shown in Figure 1(b), captions annotated by EDC contain the most comprehensive information among all methods. Figure 2 quantitatively demonstrates that the captions generated by EDC provide greatest benefit to LMMs.
We applied our EDC to annotate a large-scale dataset of 1.1 million images, consisting of 1 million diverse images (EDC-1M) and 118K real-world scene images (EDC-118K). Experiments were conducted using both LLaVA-v1.5 and LLaVA-NeXT models. The results show that the highly detailed captions generated by EDC significantly enhance the perceptual capabilities of large multimodal models (LMMs), improving visual-language alignment. The outstanding performance of both LLaVA-v1.5 and LLaVA-NeXT across 14 benchmarks further underscores the effectiveness of our approach, confirming that the generated image captions are of exceptional detail and quality.
Figure 3: The EDC pipeline first utilizes various visual specialists to extract both Object and Relation attributes. Then, it uses an LLM to integrate the object attributes into detailed region captions, followed by combining the region captions with relational attributes to generate a comprehensive image caption.
@@ 2 Related Work
2.1 Large multi-modality models
Recently, there have been significant advancements in the development of large multi-modality models (LMMs), particularly in aligning and integrating pretrained vision encoders with pretrained language models to enhance their reasoning and perception capabilities [64, 67, 59]. A key line of research focuses on designing effective bridging mechanisms between these two modalities. For instance, Flamingo [2] inserts new gated cross-attention layers between existing pretrained and frozen LM layers, that bridges powerful pretrained vision-only and language-only models. BLIP-2 [28] employs a lightweight querying transformer that facilitates interaction between vision and language representations with minimal additional computational overhead. Similarly, Qwen-VL [4] adopts an approach akin to the querying transformer, while LLaVA [36] utilizes a simple projection module to map visual embeddings to the language model’s input space. The Llama3 Herd of Models [15] extends this paradigm by incorporating an adapter mechanism akin to those in LLaVA [36] and BLIP-2 [28], further optimizing the connection between visual and language components. In contrast, Some works [74] adopts a token-based fusion strategy, directly concatenating multimodal tokens—including both image and text tokens—before processing them within a unified transformer architecture. Emu [58], on the other hand, employs a causal transformer to transform image encodings into discrete token representations, improving their compatibility with autoregressive language models.
Beyond bridging mechanisms, another research has focused on improving the efficiency and effectiveness of visual encoding architectures. One promising approach involves partitioning images into large blocks, allowing for the extraction of fine-grained features while maintaining computational efficiency. Notable works leveraging this strategy include Monkey [31], Qwen2-VL [63], and LLaVA-NeXT [27], all of which aim to enhance the granularity of visual representations while optimizing model efficiency.
In addition to architectural innovations, significant efforts have been made to improve LMMs through large-scale high-quality pretraining and fine-tuning data. High-quality and diverse datasets play a crucial role in shaping the capabilities of these models, as evidenced by works such as Qwen-VL [4], Kosmos [45], Otter [25], Shikra [6], MiniGPT [75], and CogVLM [64]. Despite the impressive progress enabled by these models, relatively few studies have systematically explored the large-scale collection and curation of high-quality image-text datasets—an essential factor in further advancing LMM capabilities. The development of improved data collection pipelines, including strategies for filtering, augmentation, and quality assessment, remains a critical open research direction for the field.
@@ 2.1 Large multi-modality models
lightgray black Depth Depth & Detection model Average the depth map values within the bounding box region to obtain the depth information.
lightgray black Emotion Emotion model If the detected region is labeled as ”person”, an emotion model is used to extract an emotion label.
lightgray black OCR OCR Model Using an OCR model to extract the text content and bounding box from the region.
lightgray black Animal Fine Grained model A fine-grained recognition model to identify specific species of the animal.
lightgray            black Plants A fine-grained recognition model to identify specific species of the plants.
lightgray            black Aircrafts A fine-grained recognition model to identify specific model of the aircraft.
lightgray            black Logo A fine-grained recognition model to identify logos in the region.
lightgray            black Landmark A fine-grained recognition model to identify landmarks within the region.
lightgray            black Food A fine-grained recognition model to identify specific species of the food.
lightgray            black Celebrity Using a fine-grained recognition model to identify celebrity within the region.
@@
lightgray black Count Detection model Counting the number of all objects in the image based on the detection results.
lightgray black 2D Absolute Location Detection model
@@
lightgray black 2D Relative Location Detection model
@@
lightgray black 3D Relative Location Detection & Depth model
@@
lightgray black
2.2 Descriptive captions
High-quality descriptive captions play a crucial role in improving the performance of LMMs by providing detailed and context-rich textual descriptions of visual content [7, 22]. Several large-scale datasets have been developed to facilitate research in image captioning and vision-language modeling [56, 61, 17]. One of the most widely used datasets, Conceptual Captions 3M (CC3M) [53], is constructed by harvesting image-alt text pairs from web data. To ensure caption quality, an automatic pipeline is employed to extract, filter, and transform candidate image-caption pairs, improving both linguistic and semantic relevance. Building upon CC3M, Conceptual Captions 12M (CC12M) [5] extends this methodology with a focus on enhancing long-tail visual recognition. CC12M significantly expands the dataset’s diversity by incorporating a broader range of objects, scenes, and rare visual concepts, thereby improving the representation of real-world imagery. In addition to automatically curated datasets, manually annotated caption datasets have been developed to ensure greater accuracy and contextual relevance. SBU Captions[44] is a large-scale dataset sourced from Flickr [47], where user-written descriptions provide natural language captions for images. These captions capture a diverse range of visual content, including both common and highly specific scenes. Similarly, COCO-Captions[9] extends the well-known COCO dataset by incorporating human-annotated image descriptions collected through Amazon Mechanical Turk (AMT). While manually curated captions tend to be more accurate and aligned with human perception, they often lack finer-grained details and contextual depth, limiting their usefulness for training models that require richer semantic understanding.
Toward this end, recent methods have leveraged LMMs to generate more detailed image captions automatically. ShareGPT4V[7] and DenseFusion[29] represent notable examples of this approach, utilizing large-scale vision-language models to generate captions with enhanced descriptive richness. These methods rely on advanced language models such as GPT-4V[7] to enrich captions with finer details that may be overlooked by human annotators. However, LMM-based caption generation still faces notable challenges, particularly regarding the accuracy and reliability of the generated descriptions. Issues such as over-simplification, hallucination, and misinterpretation of image content remain prevalent, potentially reducing caption quality and trustworthiness[46, 24]. In this paper, we propose EDC pipeline, which integrates a diverse set of visual expert models specialized in different aspects of image understanding, including object recognition, scene parsing, and fine-grained attribute detection. By leveraging multiple visual experts, our method enhances the accuracy and contextual relevance of generated captions, reducing hallucinations common in LMM-based captioning. This approach achieves a better balance between detail richness and factual correctness, advancing the state-of-the-art in descriptive caption generation.
@@ 3 Approach
As shown in Figure 3, EDC leverage existing visual specialists to extract visual properties for improving descriptive captions. We explore two kinds of properties: object-level attributes and object-relations. Our approach consists of: object localization using a SoTA open-world object detection specialist, object property extraction using various specialists, and object relation extraction between objects.
@@ 3.1 Object Attributes
Attribute extraction by visual specialists. Table 2.1 presents all the instance-level attributes involved in EDC, along with their extraction processes. The attributes currently include three main parts. (1) Fine-grained object category attributes. This is rarely explored in current multimodal models [7]. In EDC, we incorporate fine-grained details by introducing various specialized models, covering categories such as animals, plants, food, logo, aircraft, landmarks, and celebrities. The animal and plant attributes contain the fine-grained category of $891k$ and $427k$ species of animals and plants, respectively. Each species within this category possesses unique characteristics and behaviors, making it a rich and varied classification. The food attributes include a variety of food types commonly found in daily life, while logos are graphic symbols that serve as visual identifiers for conveying messages. Landmarks are the famous tourist attractions that hold cultural, historical, or geographical significance. Aircraft refers to the machines designed for flight, including airplanes, helicopters, drones, and other aerial vehicles. Celebrities are individuals widely recognized in public life, entertainment, sports, or other fields. All this specific and fine-grained information could be regarded as the external world knowledge, aligning the textual content in the image with basic human cognition. (2) Low-level and emotion attributes. It includes emotion, depth, and size. (3) OCR. It is one of the important attributes for multimodal models. The specific visual expert models we use will be presented in the Appendix A.1.
@@ 3.2 Object Relation
EDC can extract relationships between multiple objects. We consider three categories: the interactions between humans and objects, the 2D as well as 3D relative positional relationships among different objects, and the object counting information. Table 2.1 presents the relation attributes and their extraction process.
@@
Attributes InternVL2 LLaVA-NeXT EDC
@@
3.3 Captioning with Attributes
Region captioning. We use a large language model (Qwen2-72B-AWQ) to integrate the object attributes with the caption from a large multimodal model (InternVL2-26B). We design a set of structured prompts that guide LLM to generate fine-grained descriptions by integrating visual details such as category, color, texture, spatial location, and human-object interactions into coherent natural language. For example, when the fine-grained model for “animals” predicts a specific class label like $\{animal\_name\}$, we use the following prompt to guide the LLM in incorporating it into the caption:
$\{cat\_name\}$ exists in the region and $\{animal\_name\}$ is a subclass of $\{cat\_name\}$; use $\{animal\_name\}$ in the caption; otherwise, do not mention $\{animal\_name\}$
Here, $\{cat\_name\}$ denotes the coarse-grained label predicted by the detection model. The complete prompt engineering strategy is detailed in the Appendix A.3.
Table 3: Comparison of Different Image Captioning Annotation Methods. The red and blue colors respectively represent the optimal and suboptimal results on each benchmark.
   Annotation Methods    GQA    ScienceQA    MMBench    $M$-Vet    SEEDI    SEED-Bench
   LLaVA-v1.5-7B
   +ShareGPT4V [7]    63.3    68.4    68.8    37.6    69.7    61.9
   +DenseFusion [29]    64.0    69.3    69.2    37.8    70.1    62.3
   +DCI [62]    63.1    69.2    64.3    30.2    66.7    61.2
   +DOCCI [43]    63.2    69.4    65.0    32.3    66.6    61.1
   +IT [46]    62.9    68.0    65.5    33.1    67.2    61.7
   +EDC    64.2    71.0    69.2    38.2    70.3    64.3
   LLaVA-NeXT-7B
   +Recap [26]    65.0    70.5    67.2    37.2    -    64.5
   +DCI [62]    64.7    70.7    68.0    38.0    70.6    64.1
   +DOCCI [43]    64.8    71.1    68.5    36.7    70.8    64.5
   +IT [46]    64.9    71.3    68.6    38.1    71.4    65.4
   +EDC    65.2    71.2    69.3    40.1    72.2    65.7
Image captioning. We combine object relation attributes, region location information (for object grounding), and reference captions to generate improved image captions. We use a LLM to integrate object-level attributes through prompt engineering. For example, to incorporate the 3D spatial relationship between two objects, we design prompts such as:
“Relative to the camera, the $\{category\_0\}$ in $\{bbox\_0\}$ is $\{3d\_relation\}$ the $\{category\_1\}$ in $\{bbox\_1\}$.”
Here, $\{3d\_relation\}$ captures spatial cues such as “in front of” or “behind of”, which are computed based on the depth difference between the two objects. We use this mechanism to update the image caption with 3D relative localization information. The complete set of prompt templates is provided in the Appendix A.3.
3.4 Analysis
Dataset Description. We leverages EDC to generate more detailed annotations for publicly available image datasets, resulting in two enhanced datasets: (1) EDC-1M, which provides dense annotations for 1 million diverse images from Densefusion [29], covering a wide range of objects and scenes, and (2) EDC-118K, which includes refined annotations for 118,000 complex scene images from the COCO dataset [33]. We provide a detailed analysis of the EDC-annotated image captions in the Appendix B.
Attribute richness. We randomly select 100 images from the EDC-118K and invite five independent evaluators to assess the captions for each image. The evaluators analyzed and recorded the occurrence of the attributes within the captions, such as spatial relations, HOI, and OCR (optical character recognition) details.
Comparison. We conducted a comparison of human annotations, internVL2-26B, LLaVA-NeXT-34B, and EDC-annotated 118K datasets on both LLaVA-v1.5 and LLaVA-NeXT. The results as shown in Figure 2, the captions annotated by EDC significantly improve the performance of LLaVA-v1.5 and LLaVA-NeXT. Compared to human and generalist multimodal model annotations, EDC-annotated captions are richer and more detailed, offering deeper context and capturing finer nuances. This enhanced descriptive quality leads to better performance in downstream tasks.
4 Experiments